OpenStudy (anonymous):

Simplify the expression: 18+3(9f-13)

OpenStudy (welshfella):

distribut the 3 over the contents of the bracket = 18 +3*9f +3*-13 can you continue?

OpenStudy (anonymous):

okay thanks I'll try:)

OpenStudy (yanasidlinskiy):

That is incorrect. You're supposed to solve what's in parentheses first. Parentheses comes before EVEN using the distributive property.

OpenStudy (anonymous):

but you can not combine 9f and 13!

OpenStudy (yanasidlinskiy):

No wait. Sorry. I didn't even see that it was already simplified. It was an error that occurred on my end.

OpenStudy (anonymous):

I tried to distribute and this is what I got: 27f+56

OpenStudy (welshfella):

= 18 + 27f - 39 = 27f - 21